# SpikeBottleNet: Energy Efficient Spike Neural Network Partitioning for Feature Compression in Device-Edge Co-Inference Systems

### v1

@@ Abstract
The advent of intelligent mobile applications highlights the crucial demand for deploying powerful deep learning models on resource-constrained mobile devices. An effective solution in this context is the device-edge co-inference framework, which partitions a deep neural network between a mobile device and a nearby edge server. This approach requires balancing on-device computations and communication costs, often achieved through compressed intermediate feature transmission. Conventional deep neural network architectures require continuous data processing, leading to substantial energy consumption by edge devices. This motivates exploring binary, event-driven activations enabled by spiking neural networks (SNNs), known for their extremely energy efficiency. In this research, we propose a novel architecture named SpikeBottleNet, a significant improvement to the existing architecture by integrating SNNs. A key aspect of our investigation is the development of an intermediate feature compression technique specifically designed for SNNs. This technique leverages a split computing approach for SNNs to partition complex architectures, such as Spike ResNet50. By incorporating the power of SNNs within device-edge co-inference systems, experimental results demonstrate that our SpikeBottleNet achieves a significant bit compression ratio of up to 256x in the final convolutional layer while maintaining high classification accuracy with only a 2.5% reduction. Moreover, compared to the baseline BottleNet++ architecture, our framework reduces the transmitted feature size at earlier splitting points by 75%. Furthermore, in terms of the energy efficiency of edge devices, our methodology surpasses the baseline by a factor of up to 98, demonstrating significant enhancements in both efficiency and performance.
[…]
123
@@ 1 Introduction
In recent years, the progressive adoption of Deep Neural Networks (DNNs) in mobile technologies and Internet of Things (IoT) devices has catalyzed breakthrough performances in a wide array of intelligent applications [5]. However, DNN-based applications demand huge computational resources, rendering them impractical for direct implementation on resource-constrained mobile and edge devices. The most common deployment approach involves transmitting raw data to cloud or edge computing infrastructures for further processing and inference, followed by the resulting output being returned to the mobile device. While this approach is useful, it leads to considerable communication overhead, which causes an increase in both latency and energy usage. Furthermore, the computational power of IoT devices goes unused when data is entirely offloaded to servers. For applications involving personal data, such as health records, where privacy laws apply, local processing and anonymization are required, adding extra processing time. An alternative strategy is on-device inference, which entails the deployment of compressed DNNs on mobile devices. Nonetheless, this approach often results in compromised performance due to over-compression and is limited by the devices’ restricted resources, which in turn, increases the latency of on-device communication. These challenges have prompted the exploration of other solutions, with device-edge co-inference emerging as a viable option [19]. This method partitions a network, enabling a mobile device access to the leading part and an edge server to handle the rest of the computations. The key advantage of this approach lies in the reduction of feature tensor sizes in deeper layers of the network. By processing the earlier layers on a mobile device and then transmitting the resulting compressed features to the cloud, both latency and energy consumption associated with cloud data transmission can be substantially decreased.
@@
Prior research on device-edge co-inference has primarily focused on traditional deep neural networks, which rely on continuous values for activations [6][28]. Although these networks deliver high performance, they are characterized by considerable computational complexity and energy demands. In contrast, distinguished as the third generation of neural network models, the brain-inspired Spiking Neural Networks (SNNs) offer a promising alternative that is extremely energy efficient when deployed in specialized neuromorphic hardware [27][2]. The adoption of binary spike signals enables SNNs to utilize low-power sparse additions as opposed to the high-power matrix multiplication operations prevalent in traditional models. SNNs also distinguish themselves by operating on binary spikes for neural communication, thereby reducing the communication overhead. Additionally, SNNs can inherently support feature compression through sparse representations and temporal coding, minimizing the need for quantization techniques aimed at shrinking feature sizes for transmission. The combination of comparable accuracy with traditional DNN [8], reduced energy consumption [29][15][26], and a high feature compression ratio makes SNNs an optimal choice for application in device-edge co-inference contexts.
@@
This study presents a novel architecture, named SpikeBottleNet, utilizing spiking neural networks (SNNs) for image classification, emphasizing efficient feature compression within device-edge co-inference systems. The key contributions of this paper are highlighted as follows:
@@
  • The experiments on CIFAR100 Datasets demonstrate that our model achieves a compression ratio of up to 256× in the final convolutional layer while maintaining an accuracy degradation of approximately 2.5%. This performance is achieved using an architecture similar to BottleNet++, but with a significant reduction in mobile energy consumption, by a factor of up to 98.
  • A key strength of SpikeBottleNet lies in its ability to facilitate the partitioning of a SNN at earlier layers. Compared to BottleNet++, SpikeBottleNet achieves a bit reduction of up to 75% at these earlier layers. This leads up to 70× reduction in energy consumption of edge devices compared to existing compression methods.
This paper continues as organized in the following manner: Section 2 summarizes an overview of related research in the field. Section 3 delves into the essential background required for comprehending the proposed work, followed by a detailed explanation of the SpikeBottleNet architecture. Section 4 outlines the experimental setup employed for our evaluations and presents a thorough account of the results obtained. The paper concludes with Section 5, which revisits the key findings and identifies potential areas for future study.
2 Background and Related Work
@@ 2.1 Deep Spiking Residual Network
In the domain of deep learning, Residual Networks (ResNet) have emerged as a prominent architecture specifically designed to tackle the degradation issue observed in increasingly deep networks. [9]. He et al. introduced the concept of identity mapping between layers through the implementation of shortcut connections. This technique significantly facilitated the training of very deep neural networks [9]. Inspired by this breakthrough in residual learning, Zheng et al. [32] established pioneering work in SNN-oriented batch normalization (BN). Their work also demonstrated the successful training of spiking ResNet-34/50 architectures using surrogate gradients within the STBP framework [30]. Extending this progression, Fang et al. [7] introduced the spike-element-wise (SEW) ResNet, employing shortcut-based integer spike merging to extend the depth of SNNs beyond 100 layers. Furthermore, Hu et al.’s Membrane Shortcut MS-ResNet introduced a novel shortcut for neuron layer membrane potentials, facilitating the training of up to 482-layer network [13]. In our study, we chose ResNet50 for its optimal balance of depth and computational efficiency, alongside its established benchmarking
These SNN based models are well-suited for implementation on neuromorphic chips, like NorthPole [23], TrueNorth [1], Loihi [24], and ROLLS [14]. These chips contain multiple interconnected neuromorphic cores (NCs) capable of running SNNs in an energy-saving manner. Unlike traditional architectures constantly processing data, these chips only activate specific neurons when receiving input spikes. This sparse operation significantly reduces power consumption, as a large portion of the chip remains idle at any given time. Compared to DNNs, the SNNs running on neuromorphic platforms can achieve up to 1000x greater power efficiency [25]. Our proposed SpikeBottleNet architecture aligns with the spike-driven paradigm, making it particularly well-suited for implementation on neuromorphic chips.
@@ 2.2 Model Splitting
Split Computing (SC) [22] is a technique that segments a DNN model into two segments: the initial segment executes on an edge device, while the subsequent portion operates on a server. This division is designed to ease the computational workload on the edge device while reducing the costs of data transmission. This is because processing delays on an edge server are typically less significant than those encountered with local processing and data transmission delays [20]. For reducing communication cost, recent research on dividing models has introduced a bottleneck feature in vision models, which results in a smaller data size than the original input sample [19][6][28][21]. Yet, these studies have not leveraged the potential of SNNs in reducing data size through model splitting and bottleneck layer, thereby overlooking an opportunity to further optimize performance.
@@ 2.3 Feature Compression
The interaction between network splitting and feature compression presents a significant challenge in the device-edge co-inference. The ideal splitting strategy seeks to minimize both on-device communication latency and feature transmission size. Feature compression has emerged as a promising approach, with numerous studies proposing various splitting and compression techniques. For instance, Li et al. in JALAD [18] employed quantization and Huffman coding for in-layer feature map compression, reducing data volume. Another study [3] presented a deep feature compression structure comprising encoder-decoder pairs, suitable for edge-cloud collaborative inference environments. BottleNet [6] introduced a bottleneck unit with a trainable lossy compressor, leveraging compressed sensing training to mitigate accuracy degradation from lossy compression. Building upon BottleNet, BottleNet++ [28] proposed an end-to-end architecture that optimizes both feature compression and encoding by integrating deep joint source-channel coding. However, BottleNet++ neglects on-device model compression, leading to excessive on-device computations at most split points. This translates to increased inference latency and energy consumption. In a more recent work [19], a novel feature compression module employing a channel attention mechanism has been introduced. This module identifies the most critical features and employs entropy encoding for eliminating statistical redundancy within the compressed data. Notably, all aforementioned research focuses on DNNs, while the potential of SNNs remain unexplored in this domain.
Figure 1: The architecture of the SpikeBottleNet
3 Proposed Method
@@ 3.1 Overall Network Architecture
The architecture of our proposed model primarily comprises the design of the spike residual network and the bottleneck layer and positioning the bottleneck layer in different splitting points within the network to efficiently compress and transmit the cuboid intermediary feature tensors. As depicted in Figure 1, the feature compression module comprises an encoder and a decoder, which are designed as lightweight CNNs. These networks perform convolution operations on spikes produced by Leaky Integrate-and-Fire (LIF) neurons. This setup differs from that of BottleNet++ by substituting ReLU activation layer with LIF neurons. While Batch Normalization (BN) is commonly used in conventional DNNs for allowing stable convergence and enable deeper architectures, SNNs demand a tailored normalization technique due to their unique temporal dimension and activation dynamics. This is crucial to mitigate the problems of vanishing or exploding gradients and regulate firing rates. To address these specific requirements, we leverage threshold-dependent batch normalization (tdBN) as proposed by Zheng et al. [32].
@@ 3.2 Spiking Neuron Model
Neural networks are composed of neurons, which transform synaptic inputs into action potential outputs. In contrast to DNNs, which ignore temporal dynamics and focuses on spatial information transmission, SNNs employ spiking neurons. These neurons simulate the behaviour of biological plausibility through the dynamics of membrane potentials and spike-based communication. The Leaky Integrate-and-Fire (LIF) model is commonly used in SNNs because it offers an optimal balance between biological plausibility and computation complexity alongside reduced energy consumption compared to DNN neurons. Our study adopts the iterative LIF model by Wu et al. [31], applied in SNNs, which uses the Euler method to convert the LIF differential equation into a recursive formula:
@@
$V_{t,n+1}=\tau_{decay}V_{t-1,n+1}\left(1-o_{t-1,n+1}\right)+x_{t,n}$ (2)
@@
Figure 2: Architecture of the ResNet50 and highlighting its 16 residual blocks.
@@ 3.3 Threshold-dependent batch normalization
We adopt the TDBN [32] normalization method, which operate across both spatial and temporal domains. It focuses on updating the membrane potential $u_{i}^{t}$ of a neuron at time $t$, using the TDBN to normalize a mini batch of the sequential inputs $\{I_{i}^{t}=\sum_{j=1}^{n}{w_{ij}o_{j}^{t}\ |\ t=1,\ \ldots,\ T}\}$ as follows:
@@ 3.5 Bottleneck Unit
The bottleneck unit architecture significantly enhances data transmission efficiency between edge devices and the server, utilizing a feature reduction unit (encoder) and feature recovery unit (decoder). This design compresses data on edge devices before decompressing it in the cloud, restoring original dimensions. Both the encoder and decoder integrate three layers: a convolutional layer (or deconvolutional layer for the decoder), a threshold-dependent batch normalization (tdBN) layer, and an activation layer using LIF neurons. The encoder aims to compress data by reducing its spatial and channel dimensions, employing convolutional filters to adjust output channels and effectively manage spatial dimensions. In contrast, the decoder focuses on expanding the dimensions of the features to their original state, using a similar approach to the encoder but with an emphasis on dimension restoration. The tensor dimension changes in this process, starting from $(time\_step,batch\_size,c,w,h)$ and compressing to $(time\_step,batch\_size,c^{\prime},\ w^{\prime},\ h^{\prime})$. This involves narrowing the $c$ channels to $c^{\prime}$ and adjusting the spatial dimensions to $w^{\prime}$ and $h^{\prime}$, which is accomplished by choosing appropriate filter sizes for $c^{\prime}$ and adjusting convolution kernel strides to $\left\lfloor\frac{w}{w^{\prime}}\right\rfloor$ and $\left\lfloor\frac{h}{h^{\prime}}\right\rfloor$. This streamlined architecture ensures efficient data handling by minimizing transmission load while maintaining the integrity and usability of the data, making it particularly effective for applications requiring real-time data exchange between edge devices and cloud services.
@@ 3.6 Training Strategy
In this research, we introduce a novel architecture trainable in an end-to-end fashion. However, direct training of the entire architecture faces slow convergence. Therefore, we adopt a two-step training approach: initially training the SNN ResNet to achieve optimal accuracy [13]. The SNN ResNet consists of several layers and blocks, depicted in the Figure 2. The structure of ResNet, with its residual blocks, provides inherent breakpoints within the network that can be exploited in device edge co-inference scenarios. Next step, we deploy the bottleneck layer at these splitting points. In this step, we adopt the compression-aware training approach from [6]. The whole model is fine-tuned simultaneously by training and updating the compression module’s weights and the SNN’s parameters. Unlike the BottleNet++, which suggests a three-step approach by separately training the encoder and decoder layers in the second step while freezing the other parameters, we ignore that step to prevent slower training convergence that we observe in our experiments.
Our methodology involves placing the bottleneck unit after certain layers chosen from the underlying network, at up to $M$ potential splitting points, with $M$ being no greater than the network’s $N$ total blocks. For each point, we explore bottleneck layers that implement varying degrees of compression along channel or spatial dimensions. From those models that maintain an accuracy reduction within 3%, we select the model that requires the highest bit compression ratio. This selection process is replicated across all $M$ locations. Ultimately, we pick the best way to partition the network that minimizes edge device energy consumption and/or maximizes the bit compression ratio, in alignment with our optimization goals.
@@
Splitting Point RB1 RB2 RB3 RB4 RB5 RB6 RB7 RB8 RB9 RB10 RB11 RB12 RB13 RB14 RB15 RB16
BottleNet++ (B) [28] 8,192 16,384 16,384 8,192 8,192 8,192 8,192 4,096 4,096 4,096 4,096 4,096 4,096 1,024 1,024 128
SpikeBottleNet (B) 8,192 4,096 4,096 4,096 4,096 4,096 4,096 2,048 1,024 1,024 1,024 1,024 2,048 256 512 64
Bit Reduction (%) 0 75 75 50 50 50 50 50 75 75 75 75 50 75 50 50
Accuracy Drop (%) 3.45 2.33 2.13 1.40 2.94 3.70 2.34 3.63 3.90 3.13 2.64 2.85 2.61 3.76 2.53 2.52
Time Step 4 2 2 2 2 2 2 4 2 2 2 2 4 2 4 4
In the experiment, the model consistently reached the 3% accuracy degradation threshold across these splitting points. However, to investigate the trade-off between accuracy drop and compression ratio, we subsequently evaluated the model’s performance when tolerating an accuracy drop of 4%.
Table 1: Comparison of the impact of splitting points on the compressed feature size of BottleNet++ and SpikeBottleNet. Additionally, the table presents the model accuracy and bit reduction achieved by SpikeBottleNet and the Time Step used at different splitting points.

### v2

# SpikeBottleNet: Spike-Driven Feature Compression Architecture for Edge-Cloud Co-Inference

@@ Abstract
Edge-cloud co-inference enables efficient deep neural network (DNN) deployment by splitting the architecture between an edge device and cloud server, crucial for resource-constraint edge devices. This approach requires balancing on-device computations and communication costs, often achieved through compressed intermediate feature transmission. Conventional DNN architectures require continuous data processing and floating point activations, leading to considerable energy consumption and increased feature sizes, thus raising transmission costs. This challenge motivates exploring binary, event-driven activations using spiking neural networks (SNNs), known for their extreme energy efficiency. In this research, we propose SpikeBottleNet, a novel architecture for edge-cloud co-inference systems that integrates a spiking neuron model to significantly reduce energy consumption on edge devices. A key innovation of our study is an intermediate feature compression technique tailored for SNNs for efficient feature transmission. This technique leverages a split computing approach to strategically place encoder-decoder bottleneck units within complex deep architectures like ResNet and MobileNet. Experimental results demonstrate that SpikeBottleNet achieves up to 256x bit compression in the final convolutional layer of ResNet, with minimal accuracy loss (0.16%). Additionally, our approach enhances edge device energy efficiency by up to 144x compared to the baseline BottleNet, making it ideal for resource-limited edge devices.
@@ 1 Introduction
Recently, adopting Deep Neural Networks (DNNs) in mobile and Internet of Things (IoT) devices has driven breakthroughs across a wide array of intelligent applications [1]. However, DNN-based applications require substantial computational resources, and their direct implementation incurs significant power and computation costs. The most common deployment approach involves transmitting raw data to a cloud service for inference. While this approach is useful, it leads to considerable communication overhead and latency. Furthermore, the computational power of IoT devices goes unused when data is entirely offloaded to servers. For applications involving personal data, such as health records, local processing and anonymization are required.
@@
An alternative strategy is on-device inference, which entails the deployment of compressed DNNs on mobile devices. Nonetheless, this approach often results in compromised performance due to over-compression. These challenges have prompted the exploration of other solutions, with device-edge co-inference emerging as a viable option [2]. This method partitions a network, enabling a mobile device access to the leading part and an edge server to handle the rest of the computations. By processing the earlier layers on a mobile device and then transmitting the compressed features to the cloud, both latency and energy consumption associated with cloud data transmission can be substantially decreased.
@@
Prior research on device-edge co-inference has primarily focused on traditional deep neural networks, which rely on continuous values for activations [3][4]. Although these networks deliver high performance, they are characterized by considerable computational complexity and energy demands. In contrast, distinguished as the third generation of neural network models, the brain-inspired Spiking Neural Networks (SNNs) offer a promising alternative that is extremely energy efficient when deployed in specialized neuromorphic hardware [5][6]. The adoption of binary spike signals enables SNNs to utilize low-power sparse additions as opposed to the high-power matrix multiplication operations prevalent in traditional models.
SNNs can inherently support feature compression through sparse representations and temporal coding, thereby reducing the communication overhead. The combination of comparable accuracy with traditional DNN [7], reduced energy consumption [8][9][10], and a high feature compression ratio makes SNNs an optimal choice for application in device-edge co-inference contexts.
@@
This study introduces SpikeBottleNet, an SNN-based architecture for energy-efficient feature compression in device-edge co-inference image classification. The key contributions of this paper are highlighted as follows:
@@
  • The experiments on CIFAR100 datasets demonstrate that SpikeBottleNet reduces the ResNet50 and MobileNetV1 intermediate feature sizes by factors of 256x and 174x, respectively, in the final convolutional layer. The model experiences only minimal accuracy degradation of approximately 0.16% and 0.72% respectively.
  • Our model’s performance is achieved using an architecture similar to DNN based BottleNet[4], but with a significant reduction in mobile energy consumption, by a factor of 12x on traditional hardware and up to 144x on neuromorphic hardware.
The paper proceeds as follows: Section 2 provides a review of related research. Section 3 covers the necessary background and details the SpikeBottleNet architecture. Section 4 describes the experimental setup and presents the results. Finally, Section 5 summarizes key findings and suggests directions for future research.
2 Background
@@ 2.1 Deep Spiking Residual Network
Residual Networks (ResNet) address the degradation problem in deep learning, particularly in very deep networks. [11]. He et al. introduced identity mappings between layers through shortcut connections. This technique significantly facilitated the training of very deep neural networks. Inspired by this breakthrough, Zheng et al. [12] established pioneering work in SNN-oriented batch normalization (BN) and demonstrated the successful training of spiking ResNet-34/50 architectures using surrogate gradients within the STBP framework [13]. Extending this progression, Fang et al. [14] introduced the spike-element-wise (SEW) ResNet. Furthermore, Hu et al.’s Membrane Shortcut MS-ResNet introduced a novel shortcut for neuron layer membrane potentials, facilitating the training of up to 482-layer network [15]. Beyond ResNet, architectures like MobileNet have been employed in spiking neural networks for energy-efficient, competitive object detection in real-world automotive event data [16]. In our study, we chose ResNet50 for its optimal balance of depth and established benchmarks, alongside a lightweight MobileNetV1, to showcase SpikeBottleNet’s general applicability.
SNN based models are well-suited for neuromorphic chips, like NorthPole [17], TrueNorth [18], Loihi [19], and ROLLS [20]. Unlike traditional architectures constantly processing data, these chips only activate specific neurons when receiving input spikes. This sparse operation significantly reduces power consumption, as a large portion of the chip remains idle. Compared to DNNs, the SNNs running on neuromorphic platforms can achieve up to 1000x greater power efficiency [21]. Our proposed SpikeBottleNet architecture aligns with the spike-driven paradigm, making it well-suited for implementation on neuromorphic chips.
@@ 2.2 Model Splitting
Split Computing (SC) is a technique that segments a DNN model into two parts: the initial segment executes on an edge device, while the subsequent portion operates on a server [22]. This division reduces computational load on the edge device and lowers data transmission costs. Processing delays on an edge server are typically less significant than those encountered with local processing and data transmission delays [23]. Recent research on dividing models has introduced a bottleneck feature in vision models, which results in a smaller data size than the original input sample [2][3][4][24]. However, these studies have not leveraged the potential of SNNs in reducing data size through model splitting and bottleneck layer, thereby overlooking an opportunity to further optimize performance.
@@ 2.3 Feature Compression
The interaction between network splitting and feature compression presents a significant challenge in the device-edge co-inference. The ideal splitting strategy seeks to minimize both on-device communication latency and feature transmission size. Various studies have explored splitting and compression techniques, such as JALAD’s use of quantization and Huffman coding for feature compression [25], and a deep feature compression structure with encoder-decoder pairs for edge-cloud collaborative inference [26]. BottleNet [3] introduced a bottleneck unit with a trainable lossy compressor, and BottleNet++ optimized feature compression with joint source-channel coding [4]. However, BottleNet++ neglects on-device model compression, leading to excessive on-device computations at most split points. Recent work introduced a compression module using channel attention and entropy encoding [2]. Existing research has focused only on DNNs, overlooking SNNs, which naturally enable sparse data representation and lower communication overhead—ideal for efficient feature compression
Figure 1: Overview of the SpikeBottleNet architecture showing a spiking network with feature compression using an encoder-decoder module. The network partitions intermediate spike tensors at specific layers to optimize communication and energy efficiency in device-edge co-inference.
3 Methodology
@@ 3.1 Overall Network Architecture
The architecture of our proposed model primarily comprises the spike residual network and the bottleneck layer positioned at different network split points to efficiently compress and transmit the intermediary feature tensors. As depicted in Figure 1, the feature compression module comprises lightweight CNN-based encoder and a decoder. These layers perform convolution operations on spikes produced by Leaky Integrate-and-Fire (LIF) neurons. This setup differs from that of BottleNet by substituting ReLU activation layer with LIF neurons for spike-based processing. Due to the temporal dynamic of SNNs, we use threshold-dependent batch normalization (tdBN) as suggested by Zheng et al. [12] to avoid gradient issues and regulate firing rates.
@@
In contrast, the SNN MobileNet architecture leverages Inverted Residual Blocks (IRBs) which consists of depthwise separable convolution blocks (DSCs) to reduce computational complexity. Each IRB performs a depthwise convolution followed by a pointwise convolution. The architecture begins with a standard convolution and uses depthwise separable convolutions in subsequent layers. Similar to ResNet, final stage includes average pooling and a fully connected layer for classification. LIF neurons are placed after each RB in ResNet and IRB in MobileNet, as splitting points are chosen after these blocks to ensure they generate spikes for efficient feature compression and transmission.
@@ 3.2 Spiking Neuron Model
Spiking neural networks are composed of neurons, which transform synaptic inputs into action potential outputs. In contrast to DNNs, which ignore temporal dynamics and focuses on spatial information transmission, SNNs employ spiking neurons. These neurons simulate the behaviour of biological plausibility through the dynamics of membrane potentials and spike-based communication. The Leaky Integrate-and-Fire (LIF) model is commonly used in SNNs because it offers an optimal balance between biological plausibility and computation complexity alongside reduced energy consumption compared to DNN neurons. Our study adopts the iterative LIF model by Wu et al., applied in SNNs, which uses the Euler method to convert the LIF differential equation into a recursive formula [27]:
@@
$V_{t,n+1}=\tau_{decay}V_{t-1,n+1}\left(1-o_{t-1,n+1}\right)+x_{t,n}\$ (2)
@@ 3.3 Threshold-dependent batch normalization
SNNs demand a tailored normalization technique due to their unique temporal dimension and activation dynamics. We adopt the TDBN [12] normalization method, which operate across both spatial and temporal domains. It focuses on updating the membrane potential $u_{i}^{t}$ of a neuron at time $t$, using the TDBN to normalize a mini batch of the sequential inputs $\{I_{i}^{t}=\sum_{j=1}^{n}{w_{ij}o_{j}^{t}\ |\ t=1,\ \ldots,\ T}\}$ as follows:
@@ 3.5 Bottleneck Unit
The bottleneck unit architecture significantly enhances data transmission efficiency between edge devices and the server, utilizing a feature reduction unit (encoder) and feature recovery unit (decoder). Both the encoder and decoder integrate three layers: a convolutional layer (or deconvolutional layer for the decoder), a threshold-dependent batch normalization (tdBN) layer, and an activation layer using LIF neurons. The encoder reduces the spatial and channel dimensions. In contrast, the decoder restores the feature dimensions. The tensor dimension changes in this process, starting from $(time\_step,batch\_size,c,w,h)$ and compressing to $(time\_step,batch\_size,c^{\prime},\ w^{\prime},\ h^{\prime})$. This involves narrowing the $c$ channels to $c^{\prime}$ and adjusting the spatial dimensions to $w^{\prime}$ and $h^{\prime}$, which is accomplished by choosing appropriate filter sizes for $c^{\prime}$ and adjusting convolution kernel strides to $\left\lfloor\frac{w}{w^{\prime}}\right\rfloor$ and $\left\lfloor\frac{h}{h^{\prime}}\right\rfloor$. This streamlined architecture minimizes transmission load while maintaining the integrity and usability of the data, making it particularly effective for applications requiring real-time data exchange between edge devices and cloud services.
@@ 3.6 Training Strategy
In this research, we introduce a novel architecture trainable in an end-to-end fashion. However, direct training of the entire architecture faces slow convergence. Therefore, we adopt a two-step training approach: initially training the SNNs to achieve optimal accuracy [28]. The block structures of ResNet or MobileNet provides inherent breakpoints within the network that can be exploited in device edge co-inference scenarios. Next step, we deploy the bottleneck layer at these splitting points. In this step, we adopt the compression-aware training approach from [3]. The whole model is fine-tuned simultaneously by training and updating the compression module’s weights and the SNN’s parameters.
Our methodology involves placing the bottleneck unit after certain blocks chosen from the underlying network, at up to $M$ potential splitting points, with $M$ being no greater than the network’s $N$ total blocks. For each point, we explore bottleneck layers that implement varying degrees of compression along channel or spatial dimensions. From those models that maintain an accuracy reduction within 2%, we select the model that requires the highest bit compression ratio. This selection process is replicated across all $M$ locations. Ultimately, we pick the best way to partition the network that minimizes edge device energy consumption and/or maximizes the bit compression ratio, in alignment with our optimization goals.